slope -3/2 ordered pair(-2,1)
Answer by nerdybill(7384)

slope -3/2 ordered pair(-2,1)
.
plug given data into "point-slope" form:
y - y1 = m(x - x1)
y - 1 = (-3/2)(x - (-2))
y - 1 = (-3/2)(x + 2)
y - 1 = (-3/2)x - 3
y = (-3/2)x - 2 (this is what they're looking for)